# Non-degenerate mixed maps and contact structures

### v1

@@ Abstract.
We study the geometry and topology of real analytic maps $\mathbb{C}^{n}\to\mathbb{C}^{k}$, where $n>k$, regarded as mixed maps, defined below. Firstly, we give two natural families of mixed isolated complete intersection singularities, called mixed ICIS, which are interesting on their own. We consider the notion of (partial) non-degeneracy for mixed maps; we prove that these define mixed ICIS and that, under suitable conditions, admit a local Milnor fibration. Then, building on previous constructions due to Oka, we obtain natural contact structures and adapted open books on a particular class of mixed links. Finally, we look at mixed links that are diffeomorphic to holomorphic ones, and we address the problem of comparing different contact structures.
@@ Introduction
Mixed maps are real analytic maps in complex variables and their conjugates. Perhaps the first time that such maps appeared in singularity theory was in the 1973 paper [1] by N. A’Campo where he constructed non-trivial examples of real analytic maps into $\mathbb{R}^{2}$ with a local Milnor fibration; this answered a question raised by J. Milnor in his classical book [21]. Some 25 years later, the articles [34] and [33], and later Seade’s book [35], opened a line of research on Milnor fibrations for real singularities, based on the use of mixed functions (though the name mixed was not used), and highlightenning the use of complex geometry in this setting. It was M. Oka in [25] who actually coined the name “mixed singularities” and in a series of remarkable papers showed that these inherit several properties from the holomorphic context and are objects of great interest from the topological point of view. Several authors have contributed to making this a rich and interesting theory, see for instance [8], [18], [25], [26], [27], [12], [31], to cite a few. For a general account, we refer the reader to [29].
@@
In this work we study mixed isolated complete intersection singularities, mixed ICIS for short. We give in the first part two constructions of such maps. The first of these we call Siegel maps, as these spring by considering the space of Siegel leaves of a generic linear action of $\mathbb{C}^{k}$ in $\mathbb{C}^{n}$ in the Siegel domain. This is interesting on its own. The regular part of the mixed ICIS has a canonical complex structure induced by the linear action, and it admits a $\mathbb{C}^{*}$-action with compact quotient. This gives rise to an important class of complex manifolds known as LVM-manifolds, a type of moment-angle manifolds of importance in algebraic topology and mathematical physics.
@@
We next look at the notion of (strong) non-degeneracy and partial non-degeneracy for mixed maps. This extends the classical notion introduced by Khovanskii in [16] for holomorphic maps, and the partial non-degeneracy for functions considered by Mondal in [22] also in the holomorphic setting, and by Bode and Quiceno in [3] for mixed functions in two variables. We determine some general properties and prove that partial non-degeneracy implies the mixed ICIS property; we also show that under additional conditions, strong non-degeneracy yields the existence of a Milnor fibration on the sphere. This gives a second class of examples provided by a construction introduced by Oka in the case of mixed hypersurfaces called mixed coverings. As a byproduct, we give the mixed version of Hamm’s complete intersections [14].
@@
On the other hand, contact structures appear in singularity theory after the work [37] of Varchenko, who proved that links of complex analytic varieties with an isolated singularity are contact submanifolds of the sphere endowed with a canonical structure. Additionally, using the spherical Milnor fibrations, Giroux proved in [11] that they are closely related to open book decompositions. Since then, Oka and others authors have studied contact structures associated to real and complex singularities.
We extend previous constructions due to Oka in [28], giving sufficient conditions on mixed links associated with certain classes of non-degenerate mixed maps, to admit a canonical contact structure. In particular, we generalize Oka’s results, in which pullbacks of partially non-degenerate holomorphic maps are used to construct mixed maps through homogeneous mixed coverings. We show that the associated links are also contact submanifolds of the sphere, and we prove the existence of adapted open books, based on the works [6] and [5]. The argument considers a suitable change of the defining contact form developed in [6] and [28], and a formulation of a condition associated with non-isolated complex singularities in [6]. Furthermore, it happens that some classes of links defined by mixed functions are diffeomorphic to holomorphic ones. In the last part, we discuss the relation between the natural contact structure of the mixed link (when this exists) and the one induced from the holomorphic case.
@@ 1.1. Definitions
A critical point of a mixed map germ is a point for which the rank of the Jacobian matrix is not maximal and we call it a mixed singularity. The set of critical points of a map $F$ is denoted by $\Sigma_{F}$ and its zero set $F^{-1}(0)$ by $V_{F}$. We have the following characterization of mixed singularities stated in [31, Proposition 4].
@@
(1.2) $\alpha_{1}{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{1}(a)+\dots+\alpha_{k}{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{k}(a)={}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}{}\mkern 3.0mu\overline{\mkern-3.0muDf_{1}}(a)+\dots+{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}{}\mkern 3.0mu\overline{\mkern-3.0muDf_{k}}(a).$
This generalizes [25, Proposition 1], which states the following.
@@ 1.2. Linear actions on $\mathbb{C}^{n}$
These are the twisted Pham-Brieskorn polynomials from [34] and [33], where it is proved that if the $a_{i}$ are all $\geq 2$, then these have a unique critical point at 0 and they have a Milnor fibration. This was the birth of the theory of mixed functions. In fact these singularities have a canonical action of $\mathbb{R}\times\mathbb{S}^{1}$, later called a polar action [8]. When the permutation $\sigma$ is the identity, these were called mixed Pham-Brieskorn polynomials in [25]. Of course one may consider now several of these equations and ask under which conditions the resulting map is a mixed complete intersection. A particular case corresponds to the mixed Hamm complete intersections, discussed in Subsection 2.2.
@@ 1.3. Mixed coverings
We now introduce a method due to Oka in [28] that allows constructing mixed maps out from holomorphic ones. He used this to construct interesting mixed functions and mixed hypersurfaces. This method also works for complete intersections.
@@
If there exist positive integers $a\neq b$ such that $a_{i}=a$ and $b_{i}=b$ for all $i$, then $\phi_{\textbf{a},\textbf{b}}$ is called a homogeneous mixed covering and denoted by $\phi_{a,b}$. Observe that $\phi:(\mathbb{C}^{*n},0)\longrightarrow(\mathbb{C}^{*n},0)$ is a diffeomorphism. Notice yet the similarity of this construction with the mixed maps (1.3).
@@ 1.4. Algebraic ICIS
Let $\Psi$ be a Siegel complete intersection map determined by an admissible configuration $\Lambda=(\Lambda_{1},\dots,\Lambda_{n})$. We claim that it is not an algebraic ICIS. For this, consider $\Psi:(\mathbb{R}^{2n},0)\longrightarrow(\mathbb{R}^{2k},0)$ and its complexification $\Psi_{\mathbb{C}}:(\mathbb{C}^{2n},0)\longrightarrow(\mathbb{C}^{2k},0)$. Its coordinate functions are:
$\operatorname{Re}\Psi^{i}_{\mathbb{C}}=\sum_{j=1}^{n}\operatorname{Re}\lambda_{ij}\left(\xi_{i}^{1}+\xi_{i}^{2}\right)\quad\hbox{and}\quad\operatorname{Im}\Psi^{i}_{\mathbb{C}}=\sum_{j=1}^{n}\operatorname{Im}\lambda_{ij}\left(\xi_{i}^{1}+\xi_{i}^{2}\right)\,$,
where $\xi_{i}^{1},\xi_{i}^{2}$ are complex variables. Take the union $L$ of the complex lines $L_{i}=\{\xi_{i}^{1}=\pm{\bf i}\xi_{i}^{2}\}$. At a point $p\in L$, $\Psi_{\mathbb{C}}(p)=0$ and its Jacobian matrix has pairwise linearly dependent columns. In other words, $L\subset\Psi_{\mathbb{C}}^{-1}(0)\cap\Sigma_{\Psi_{\mathbb{C}}}$ and $\Psi$ is not an algebraic ICIS because this property does not hold for its complexification.
@@
    A mixed monomial $f_{\mu,\nu}=\lambda_{\mu,\nu}z^{\mu}{}\mkern 3.0mu\overline{\mkern-3.0muz}^{\nu}$ is called purely mixed with respect to $z_{i}$ if $\mu_{i},\nu_{i}\geq 1$ and $\mu_{i}+\nu_{j}\geq 3$.
@@
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map such that $f^{i}$ is purely mixed for every $i=1,\dots,k$. Then $F$ is not an algebraic mixed ICIS.
@@
for some index $j$ and a mixed function $h_{\mu,\nu}$. Similar expressions hold for $\operatorname{Im}f_{\mu,\nu}$ and the derivatives with respect to $\mkern 3.0mu\overline{\mkern-3.0muz}_{i}$. Applying it to every monomial, one can see that $f_{\mathbb{C}}$ and its derivative vanish on the set $L$ consisting of the union of the zero sets correspondent to the complexification of $\|z_{i}\|^{2}$. Therefore, since each coordinate function $f^{i}$ of $F$ is purely mixed, $L\subset F_{\mathbb{C}}^{-1}(0)\cap\Sigma_{F_{\mathbb{C}}}$ and the result follows. ∎
As a consequence, mixed maps constructed as the pullback of mixed coverings are never algebraic mixed ICIS. This illustrates the substantial differences between the mixed and holomorphic settings (see Subsection 3.2).
@@ 2. Non-degeneracy
In this section we extend the notions of non-degeneracy for mixed maps. A mixed function germ is denoted by $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ and a mixed map germ by $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$. From now on we assume that $V_{f}$ and $V_{F}$ have positive dimension in $\mathbb{C}^{n}$.
In [26], M. Oka introduced the notion of the Newton polyhedra of a mixed function germ $f$, denoted by $\Gamma^{+}(f)$, and defined the condition of (strong) non-degeneracy. If $f$ is holomorphic, these two notions coincide with the classical definitions due to Kushnirenko in [17].
Let us fix some notations. Let $I\subset\{1,\dots,n\}$ be a non-empty subset. We define $\mathbb{C}^{I}=\{z\in\mathbb{C}^{n}:z_{i}=0\;\forall\;i\notin I\}$. We denote the face function of $f$ with respect to a vector $P$ of positive integers as $f_{P}$ and its restriction to $\mathbb{C}^{I}$ by $f^{I}$. A Newton polyhedra $\Gamma^{+}$ is called convenient if it intersects each non-empty subspace $\mathbb{R}^{I}$. P. Mondal in [22] introduced the following condition, whose main advantage is that it avoids the convenience hypothesis on the Newton polyhedron in some applications.
@@
A holomorphic function $f(z)$ is called partially non-degenerate if for every non-empty subset $I\subset\{1,\dots,n\}$ and every vector $P$ of positive integers, the vector field $(Df)^{I}_{P}(a)\neq 0$ for all $a\in\mathbb{C}^{*n}$.
For holomorphic function germs $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ these notions are related as follows. If $f$ is non-degenerate and convenient, then $f$ is partially non-degenerate by [22, Proposition X1.7]. Moreover, if $f$ is partially non-degenerate, then it has an isolated singularity at the origin as proved in [22, Theorem X1.3]. However, there are examples in which $f$ is degenerate but partially non-degenerate, see [22, Example XI.6] and Example 2.8. More generally, these implications depend on the fact that $(Df)^{I}_{P}$ and $(Df_{P})^{I}$ do not coincide in general for a subset $I\subset\{1,\dots,n\}$ and $P$ a vector of positive integers.
@@
Let $f(z)$ be a partially non-degenerate holomorphic function. There does not exist a nonzero real analytic curve $w(t):(0,1]\longrightarrow\mathbb{C}^{n}\setminus\{0\}$ such that $\lim_{t\to 0}w(t)=0$ and $f_{z_{j}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)\equiv 0$ for all $j=1,\dots,n$.
@@
Suppose the existence of such a curve. First, define $I=\{j:w_{j}(t)\not\equiv 0\}$. For each $j\in I$, we write the coordinate $w_{j}$ of $w$ as $w_{j}(t)=a_{j}t^{p_{j}}+o(t)$, where $a_{j},p_{j}\neq 0$ and $o(t)$ denotes higher order terms. One can see that
$\frac{\partial f}{\partial z_{j}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)=\left(\frac{\partial f}{\partial z_{j}}\right)^{I}_{P}(a)t^{d-p_{j}}+o(t),$
where $P=(p_{1},\dots,p_{n})$ and $d$ is the weighted homogeneous degree of $f_{P}$. We conclude that $\left(Df\right)_{P}^{I}(a)=0$, which contradicts the partial non-degeneracy condition. ∎
Mixed functions with an isolated singularity at the origin can be constructed from holomorphic ones as follows.
@@
Otherwise, by the Curve Selection Lemma and Corollary 1.2, one can find real analytic curves $\lambda(t)\subset\mathbb{S}^{1}$ and $w(t)\subset\Sigma_{g}$ such that $\lim_{t\to}w(t)=0$ and
$\overline{f_{z_{j}}}(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))a_{j}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}(t)^{a_{j}-1}w_{j}^{b_{j}}=\lambda(t)b_{j}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{b_{j}-1}w_{j}(t)^{a_{j}}f_{z_{j}}\left(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\right).$
Let $I=\{j:w_{j}(t)\not\equiv 0\}$. Since $a_{j}\neq b_{j}$ and taking norms, the expression above implies that $f_{z_{j}}$ vanishes on $\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))$ for all $j\in I$. This leads to a contradiction with Lemma 2.2. ∎

### v3

@@ Abstract.
We study the geometry and topology of real analytic maps $\mathbb{C}^{n}\to\mathbb{C}^{k}$, where $n>k$, regarded as mixed maps, defined below. Firstly, we give two natural families of mixed isolated complete intersection singularities, called mixed ICIS, which are interesting on their own. We consider the notion of Newton non-degeneracy for mixed maps; we prove that these define mixed ICIS and relate this property to the Milnor sets and the existence of Milnor fibrations. Then, building on previous constructions due to Oka, we investigate natural contact structures and adapted open books on a particular class of mixed links. Finally, we look at mixed links that are diffeomorphic to holomorphic ones, and we address the problem of comparing different contact structures.
@@ Introduction
Mixed maps are real analytic maps in complex variables and their conjugates. Perhaps the first time that such maps appeared in singularity theory was in the 1973 paper [1] by N. A’Campo where he constructed non-trivial examples of real analytic maps into $\mathbb{R}^{2}$ with a local Milnor fibration; this answered a question raised by J. Milnor in his classical book [24]. Some 25 years later, the articles [36] and [38], and later Seade’s book [39], opened a line of research on Milnor fibrations for real singularities, based on the use of mixed functions (though the name mixed was not used), and highlighting the use of complex geometry in this setting. It was M. Oka in [29] who actually coined the name “mixed singularities” and, in a series of remarkable papers, showed that these inherit several properties from the holomorphic context and are objects of great interest from the topological point of view. Several authors have contributed to making this a rich and interesting theory, see for instance [11], [14], [21], [29], [30], [31], [34], to cite a few. For a general account, we refer the reader to [28].
@@
In this work, we study mixed isolated complete intersection singularities, mixed ICIS for short. In the first part, we give two constructions of such maps. The first of these we call Siegel maps, as these spring by considering the space of Siegel leaves of a generic linear action of $\mathbb{C}^{k}$ in $\mathbb{C}^{n}$ in the Siegel domain. This is interesting on its own. The regular part of the mixed ICIS has a canonical complex structure induced by the linear action, and it admits a $\mathbb{C}^{*}$-action with compact quotient. This gives rise to an important class of complex manifolds known as LVM-manifolds, a type of moment-angle manifolds of importance in algebraic topology and mathematical physics.
@@
We next look at the notion of (strong) non-degeneracy for mixed maps. This extends the classical notion introduced by Khovanskii in [19] for complex-valued holomorphic maps. We determine some general properties and relate non-degeneracy with the mixed ICIS property, Milnor sets, and the existence of Milnor fibrations. This provides a second class of examples given by a construction introduced by Oka in the case of mixed hypersurfaces, called mixed coverings. As a byproduct, we give the mixed version of Hamm’s complete intersections [16].
@@
Contact structures appear in singularity theory after Varchenko’s work [42], which showed that links of complex analytic varieties with isolated singularities are contact submanifolds of the sphere with its canonical structure. Using the spherical Milnor fibration, Giroux further established in [13] their close relation with open book decompositions. Since then, Oka and others have studied contact structures associated with real and complex singularities.
We extend Oka’s constructions in [32] and give conditions under which mixed links associated with certain non-degenerate mixed maps admit canonical contact structures. We derive an open book decomposition based on [6] and [32], using a modification of the defining contact form and a formulation in [6] related to non-isolated singularities. Finally, we examine cases in which mixed ICIS links are diffeomorphic with holomorphic ones and compare their natural contact structures.
@@ 1.1. Definitions
A critical point of a mixed map germ is a point for which the rank of the Jacobian matrix is not maximal and we call it a mixed singularity. The set of critical points of a map $F$ is denoted by $\Sigma_{F}$ and its zero set $F^{-1}(0)$ by $V_{F}$. We have the following characterization of mixed singularities stated in [34, Proposition 4].
@@
(1.2) $\alpha_{1}{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{1}(a)+\dots+\alpha_{k}{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{k}(a)=-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}{}\mkern 3.0mu\overline{\mkern-3.0muDf_{1}}(a)-\dots-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}{}\mkern 3.0mu\overline{\mkern-3.0muDf_{k}}(a).$
This generalizes [29, Proposition 1], which states the following.
@@ 1.2. Linear actions on $\mathbb{C}^{n}$
These are the twisted Pham-Brieskorn polynomials from [38] and [36], where it is proved that if the $a_{i}$ are all $\geq 2$, then these have a unique critical point at 0 and they have a Milnor fibration. This was the birth of the theory of mixed functions. In fact these singularities have a canonical action of $\mathbb{R}\times\mathbb{S}^{1}$, later called a polar action in [11]. When the permutation $\sigma$ is the identity, these were called mixed Pham-Brieskorn polynomials in [29]. Of course one may consider now several of these equations and ask under which conditions the resulting map is a mixed complete intersection. A particular case corresponds to the mixed Hamm complete intersections, discussed in Subsection 2.2.
@@ 1.3. Mixed coverings
We now introduce a method due to Oka, which appears in [32], that allows constructing mixed maps out from holomorphic ones. He used this to construct interesting mixed functions and mixed hypersurfaces. This method also works for complete intersections.
@@
If there exist positive integers $a\neq b$ such that $a_{i}=a$ and $b_{i}=b$ for all $i$, then $\phi_{\textbf{a},\textbf{b}}$ is called a homogeneous mixed covering and denoted by $\phi_{a,b}$. Observe that $\phi:\mathbb{C}^{*n}\longrightarrow\mathbb{C}^{*n}$ is a diffeomorphism on a small neighborhood of the origin. Notice yet the similarity of this construction with the mixed maps (1.3).
@@ 1.4. Algebraic ICIS
Let $\Psi$ be a Siegel complete intersection map determined by an admissible configuration $\Lambda=(\Lambda_{1},\dots,\Lambda_{n})$. We claim that it is not an algebraic ICIS. Let us consider $\Psi:(\mathbb{R}^{2n},0)\longrightarrow(\mathbb{R}^{2k},0)$ and its complexification $\Psi_{\mathbb{C}}:(\mathbb{C}^{2n},0)\longrightarrow(\mathbb{C}^{2k},0)$. Its coordinate functions are:
$\operatorname{Re}\Psi^{i}_{\mathbb{C}}=\sum_{j=1}^{n}\operatorname{Re}\lambda_{ij}\left(\xi_{1,j}^{2}+\xi_{2,j}^{2}\right)\quad\hbox{and}\quad\operatorname{Im}\Psi^{i}_{\mathbb{C}}=\sum_{j=1}^{n}\operatorname{Im}\lambda_{ij}\left(\xi_{1,j}^{2}+\xi_{2,j}^{2}\right)\,$,
where $\xi_{1,j},\xi_{2,j}$ are complex variables. Take the intersection $L\subset\mathbb{C}^{2n}$ of the complex lines $L_{i}=\{\xi_{1,i}=\pm{\bf i}\xi_{2,i}\}$. At a point $p\in L$, $\Psi_{\mathbb{C}}(p)=0$ and its Jacobian matrix has pairwise linearly dependent columns. In other words, $L\subset\Psi_{\mathbb{C}}^{-1}(0)\cap\Sigma_{\Psi_{\mathbb{C}}}$ and $\Psi$ is not an algebraic ICIS because this property does not hold for its complexification (see [43, Proposition 1.7]).
@@
    A mixed monomial $f_{\mu,\nu}=\lambda_{\mu,\nu}z^{\mu}{}\mkern 3.0mu\overline{\mkern-3.0muz}^{\nu}$ is called purely mixed with respect to $z_{i}$ if $\mu_{i},\nu_{i}\geq 2$.
@@
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map such that $f^{j}$ is purely mixed for every $j=1,\dots,k$. Then $F$ is not an algebraic mixed ICIS.
@@
for some index $j$ and a mixed function $h_{\mu,\nu}$. Similar expressions hold for $\operatorname{Im}f_{\mu,\nu}$ and the derivatives with respect to $\mkern 3.0mu\overline{\mkern-3.0muz}_{i}$. Applying it to every monomial, one can see that $f_{\mathbb{C}}$ and its derivative vanish on the set $L$ consisting of the intersection of the zero sets corresponding to the complexification of the quadratic terms $\|z_{j}\|^{2}$. Therefore, since each coordinate function $f^{i}$ of $F$ is purely mixed, $L\subset F_{\mathbb{C}}^{-1}(0)\cap\Sigma_{F_{\mathbb{C}}}$ and the result follows. ∎
As a consequence, the pullbacks by mixed coverings (see Subsection 1.3) generate examples of mixed maps which are not algebraic ICIS. This illustrates the substantial differences between mixed and holomorphic settings (see also Subsection 2.2).
@@ 2. Non-degeneracy
In this section, we extend the notions of non-degeneracy for mixed maps. A mixed function germ is denoted by $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ and a mixed map germ by $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$. From now on we assume that $V_{f}$ and $V_{F}$ have positive dimension in $\mathbb{C}^{n}$.
In [30], M. Oka introduced the notion of the Newton polyhedron of a mixed function germ $f$, denoted by $\Gamma^{+}(f)$, and defined the condition of (strong) non-degeneracy. We denote by $\Gamma(f)$ the union of the compact faces of $\Gamma^{+}(f)$, called Newton boundary. If $f$ is holomorphic, these two notions coincide with the classical definitions due to Kushnirenko in [20].
Let us fix some notations. Let $I\subset\{1,\dots,n\}$ be a non-empty subset. We define $\mathbb{C}^{I}=\{z\in\mathbb{C}^{n}:z_{i}=0\;\forall\;i\notin I\}$. We denote the face function of $f$ with respect to a vector $P$ of positive integers as $f_{P}$ and its restriction to $\mathbb{C}^{I}$ by $f^{I}$. A Newton polyhedron $\Gamma^{+}$ is called convenient if it intersects each coordinate axis. P. Mondal in [25] introduced the following condition, whose main advantage is that it avoids the convenience hypothesis on the Newton polyhedron in some applications.
@@
A holomorphic function $f(z)$ is called partially non-degenerate if for every vector $P$ of positive integers and every non-empty subset $I\subset\{1,\dots,n\}$, the gradient $(Df)^{I}]_{P}$, denoted by $\left(Df\right)_{P}^{I}$ when there is no confusion, does not vanish in $\mathbb{C}^{*n}$.
For holomorphic function germs $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ these notions are related as follows. If $f$ is non-degenerate and convenient, then $f$ is partially non-degenerate by [25, Proposition X1.7]. Moreover, if $f$ is partially non-degenerate, then it has an isolated singularity at the origin as proved in [25, Theorem X1.3]. However, there are examples in which $f$ is degenerate but partially non-degenerate, see [25, Example XI.6]. More generally, these implications depend on the fact that $(Df)^{I}_{P}$ and $(Df_{P})^{I}$ do not coincide in general for a subset $I\subset\{1,\dots,n\}$ and $P$ a vector of positive integers.
The next two results seem to be widely known in the literature in the case of the usual notion of non-degeneracy, although we did not find a proper reference. Thus, we include the proofs for convenience and future reference.
@@
Let $f(z)$ be a partially non-degenerate holomorphic function. There does not exist a nonzero real analytic curve $z(t):(0,1]\longrightarrow\mathbb{C}^{n}\setminus\{0\}$ such that $\lim_{t\to 0}z(t)=0$ and $f_{z_{j}}(z(t))\equiv 0$ for all $j=1,\dots,n$.
@@
First, define $I=\{i:z_{i}(t)\not\equiv 0\}$. For each $i\in I$, we write the coordinate $z_{i}$ of $z$ as $z_{i}(t)=a_{i}t^{p_{i}}+o(t)$, where $a_{i},p_{i}\neq 0$ and $o(t)$ denotes higher order terms. One can see that:
$f_{z_{i}}(z(t))=f_{z_{i}}^{I}(z(t))=\left(f_{z_{j}}\right)^{I}_{P}(a)t^{d_{i}}+o(t),$
where $P=(p_{1},\dots,p_{m})$ and $d_{i}$ is the weighted homogeneous degree of the face function $(f_{z_{i}}^{I})_{P}$ of $f_{z_{i}}^{I}$. We conclude that $\left(Df\right)_{P}^{I}(a)=0$, which contradicts the partial non-degeneracy condition. ∎
Mixed functions with an isolated singularity can be constructed from holomorphic ones through mixed coverings as follows.
@@
Otherwise, by the Curve Selection Lemma and Proposition 1.2, one can find real analytic curves $\lambda(t)\subset\mathbb{S}^{1}$ and $w(t)\subset\Sigma_{g}$ such that $\lim_{t\to 0}w(t)=0$ and
$\overline{f_{z_{j}}}(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))a_{j}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}(t)^{a_{j}-1}w_{j}^{b_{j}}=\lambda(t)b_{j}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{b_{j}-1}w_{j}(t)^{a_{j}}f_{z_{j}}\left(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\right),$
for every $j=1,\dots,n$. Since $a_{j}\neq b_{j}$, if we take the norms of both sides, the expression above implies that $f_{z_{j}}\left(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\right)\equiv 0$ for all $j$. This leads to a contradiction with Lemma 2.2. ∎
Remark 2.4.
In particular, if we take $c_{i}=1$ and $d_{i}=0$ for every $i=1,\dots,n$ in the mixed covering $\phi_{\textbf{c},\textbf{d}}$, we recover that partially non-degenerate holomorphic map germs have an isolated singularity at the origin.